OpenStudy (anonymous):

If I add 9 to my secret number, I'll get a number that is 90% more than the number I'll get if I subtract 9 from my secret number. What is my secret number?

OpenStudy (anonymous):

thanjks for helping me

OpenStudy (bibby):

it's a secret ;)

OpenStudy (anonymous):

help me

OpenStudy (anonymous):



OpenStudy (anonymous):

i think it is 21

OpenStudy (anonymous):

Let the number be x. \[x+9=x-9+\frac{ 90 }{ 100 } x \] find x.

OpenStudy (bibby):

you have to be patient, or try the problem yourself working backwards

OpenStudy (anonymous):

is it 21

OpenStudy (anonymous):



OpenStudy (bibby):

the way I read it is: If I add 9 to my secret number, I'll get a number that is 90% more than the number I'll get if I subtract 9 from my secret number. \(\huge x+9=0.9(x-9)\)